OpenStudy (anonymous):

graph of 2x + 3y = 12 is shifted down 3 units. What is the equation of the shifted graph? Please show me the steps

OpenStudy (anonymous):

when you shift a linear graph of the form y = mx + c, its only the y intercept value that will change, in this case, c so your original eq was : 2x +3y = 12 where 12 is the y intercept...when it says shifted down 3 units it means it has decreased by 3 units vertically... put this into the above form and you get y = (-2/3)x + 4 so youre new eq would be y = (-2/3)x + 1

OpenStudy (anonymous):

or 3y + 2x = 3

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem :)